# FED: Fast and Efficient Dataset Deduplication Framework with GPU Acceleration

### v1

@@ 4 Optimizations and GPU Acceleration
Since typical libraries do not support $sim_{sig}(S_{i},S_{j})$, we implement a custom GPU kernel. Replacing the equality operator (==absent$=$ =) with the multiplication operator ($\times$) makes the operation matrix multiplication. Thus, the GPU kernel performs $sim_{sig}(S_{i},S_{j})$ in a matrix multiplication-like manner. The kernel exploits tiling and shared memory Harris (2012) to optimize memory access patterns and enhance parallel processing performance. After comparing $c$ signature vectors pairwise to calculate the Jaccard similarity, the index information of pairs with a Jaccard similarity higher than the predefined threshold is sent to the CPU.
@@ 5 Experiment
Table 1: System configuration
CPU 2 $\times$ AMD EPYC 7502 32-Core Processor Memory 8 $\times$ 64GB DDR4 DIMM GPU 4 $\times$ NVIDIA Tesla V100 OS Ubuntu 20.04.6 (kernel 5.4.0-100) Compiler nvcc 12.4, GCC 9.4 GPU Driver 520.61.05 CUDA Version 12.4
5.1 Target System Configuration
We use a 4-node GPU cluster for our experiments. Each node is equipped with four NVIDIA Tesla V100 GPUs, with each GPU having 32GB of memory. Table  1 provides the detailed target system configuration.
Table 2: The deduplication time in seconds on a single node is compared across the CPU baseline, the GPU baseline, and FED. The CPU baseline uses 64 CPU cores. Both the GPU baseline and FED are evaluated using four GPUs.
Method Dataset Generation Comparison Union Total Speedup
CPU baseline RealNews 18,361.1 2,983.7 17.6 21,361.4 1.0
(SlimPajama Shen et al. (2023)) C4 22,748.4 3,878.8 33.1 26,660.3 1.0
GPU baseline RealNews 1,102.4 2,016.1 36.1 3,154.7 6.8
(NVIDIA NeMo Curator Jennings et al. ) C4 1,340.8 2,992.6 31.1 4,364.5 6.1
FED RealNews 10.5 354.5 1.8 366.7 58.3
C4 13.1 696.3 1.9 711.4 37.5
5.2 Comparison Baselines
As mentioned in Section 2.4, the CPU baseline is the CPU-based MinHash LSH implementation found in SlimPajama Shen et al. (2023). We identified and fixed some bugs in the implementation. The GPU baseline is the GPU-accelerated MinHash LSH implementation in NVIDIA NeMo Curator Jennings et al. .
5.3 Datasets
The datasets we used in our experiments are the RealNews dataset Zellers et al. (2019) and C4 Raffel et al. (2020). The RealNews dataset is a large English corpus of news articles from Common Crawl, and C4 is a filtered version of Common Crawl. For C4, we sampled 100GB out of the total 700GB. We choose these datasets for two reasons. They are large-scale datasets containing over 30 million documents. We select a sufficiently large dataset to understand how long the traditional MinHash LSH method takes to process it and to see FED achieves how much acceleration. It is known from a previous study Lee et al. (2022) that deduplication of the datasets we choose can improve the performance of language models.
Preprocessing the datasets.
@@ 5.3 Datasets
For MinHash LSH, we use the default settings of the CPU baseline, 128 hash functions with $b=16$ and $r=8$. Following the approach by Lee et al. Lee et al. (2022), we create shingles using five-grams and set the Jaccard similarity threshold to 0.8. The default settings of the CPU and GPU baselines are used for parameters other than the above-mentioned. As mentioned in Section 3.2.3, the number of buckets $K$ is obtained on the fly based on the number of documents $N$, such that $K=2N^{1/2}$.
@@ 5.4 Processing Speed
We measure three different types of execution time for deduplication speed comparison: generation, comparison, and union times. The generation time involves computing the MinHash signature matrices and assigning bucket IDs for each band. The comparison time involves finding documents with the same bucket ID and calculating the Jaccard similarity between them. Finally, the union time involves constructing a union graph structure for all documents and ultimately identifying the documents to be removed.
We use RealNews and C4, which contain 32M and 50M documents, respectively. The CPU baseline processing time is measured using Python’s multiprocessing module across 64 CPU cores, while both the GPU baseline and FED are measured using four V100 GPUs. The results are summarized in Table 2.
@@
For deduplication on the RealNews dataset, FED is about 58.3 times faster than the CPU baseline and about 8.6 times faster than the GPU baseline. Notably, due to the introduction of reusable hash functions, the MinHash generation phase alone achieves roughly 1,800 times faster than the CPU baseline and 110 times faster than the GPU baseline. Similarly, on the C4 dataset, FED achieves a speedup of 37.5 over the CPU baseline and about 6.1 over the GPU baseline.
@@
The improvement in the comparison stage is less pronounced than in the generation stage. As described in Section 3.1, FED performs more actual comparisons than others. While it is possible to reduce the number of comparisons by following the method used in the CPU or GPU baseline, the FED implementation prioritizes minimizing false negatives by performing more comparisons. The result shows that performing more comparisons is feasible because optimized GPU kernels enable fast computation.
@@
The processing speed for all methods is faster on the RealNews dataset than on the C4 dataset. It is attributed to the distinct characteristics of the two datasets. For both datasets, we use 100GB of data; however, RealNews consists of 32M documents with an average length of 3,335.17 characters (including spaces), whereas C4 consists of 50M documents with an average length of 2,191.15 characters (including spaces). Consequently, the higher number of documents in C4 increases the time taken for the generation and comparison phases.
Table 3: Comparing the deduplication accuracy of the MinHash LSH-based approaches and the standard MinHash algorithm. The Ratio column represents the total number of documents that have near-duplicate pairs divided by the total number of documents. Jaccard column shows the Jaccard similarity between the sets of all near duplicates by MinHash and each method.
(a) 0.1M documents
Ratio Jaccard Time
CPU MinHash 7,317 / 0.1M - 12,800.69 s
CPU baseline 7,284 / 0.1M 0.995 440.19 s
GPU baseline 7,144 / 0.1M 0.953 172.05 s
FED 7,175 / 0.1M 0.956 1.45 s
(b) 1M documents
Ratio Jaccard Time
CPU MinHash 289,229 / 1M - 9 days
CPU baseline 288,683 / 1M 0.998 1,326.03 s
GPU baseline 285,134 / 1M 0.976 522.12 s
FED 283,248 / 1M 0.966 4.41 s
@@ 5.5 Deduplication Accuracy
We use a 0.1M and 1M dataset sampled from RealNews for this experiment because the MinHash algorithm is time-consuming on large datasets. Our sampling is not entirely random to ensure many duplicated pairs even within the two smaller datasets. We first use MinHash LSH to identify similar groups, then select a proportion of each group to construct the subset.
@@
We evaluate the accuracy of deduplication methods by measuring the Jaccard similarity between the set of all near duplicates obtained using MinHash and that obtained using each method. A higher similarity indicates that the method identifies duplicate pairs more similarly to MinHash.
Table 3 compares the deduplication accuracy of each method. FED achieves a Jaccard similarity of 0.95 or higher to the CPU MinHash result. A similarity score of 0.95 or higher means that the method has identified near-duplicate documents almost identically. In fact, on the 1M dataset, out of the 283,248 documents identified as having duplicate pairs by FED, 281,355 overlapped with those identified by MinHash. This indicates that our method not only achieves faster deduplication than other methods but also identifies near-duplicate documents similar to the accurate CPU MinHash.
@@
Generation 51.78 24.86 10.47 7.31 3.58
Comparison 1,098.48 606.25 354.45 228.66 141.47
Union 0.89 1.36 1.81 1.99 1.65
Total Time 1,151.15 632.47 366.73 237.96 146.70
@@ 5.6 Scalability
Table 6 shows the results of measuring execution time when varying the number of GPUs. Note that eight or more GPUs involve the multi-node system configuration. We use the RealNews dataset, which has 32 million documents. Even with just a single GPU, the method is nearly 20 times faster than the CPU baseline, and with 16 GPUs, the deduplication process completes in less than 147 seconds. This experiment demonstrated that our method works effectively in multi-node environments, and with more available node resources, even faster processing is possible.
@@
We increase the dataset size to verify whether FED can process trillions of tokens at a practical processing speed. The results of deduplication using 16 GPUs on the RealNews and RedPajama-1T Computer (2023) datasets are shown in Table 6. Recent LLMs are typically trained with more than trillions of tokens. The results demonstrate that FED can process trillions of tokens in just 5.1 hours.
@@
The number of hash functions used in the MinHash generation step is a key parameter that could impact the deduplication performance. For example, Rae et al. (2021) use 450 hash functions. We vary the number of hash functions from 128 to 256 and 512, and Table 6 shows the results. Despite changing the number of hash functions, FED maintained a Jaccard similarity of 0.92 or higher to the MinHash results, demonstrating that FED remains robust even with variations in the number of hash functions.
@@ 6 Conclusion
As the amount of data used for training LLMs continues to grow, the time required for data preprocessing has become increasingly significant. We present a framework, called FED, which performs MinHash LSH-based deduplication efficiently on GPUs. FED significantly outperforms the CPU-based deduplication tool included in SlimPajama by up to 58.3 times and the GPU-based deduplication tool included in NVIDIA NeMo Curator by up to 8.6 times with a node of four GPUs. Deduplication of 30 billion tokens by FED takes about 147 seconds and 1.2 trillion tokens takes about five hours in a multi-node and multi-GPU environment. Similar documents identified by FED show a Jaccard similarity of 0.95 or higher with those found using the original MinHash algorithm, indicating that FED achieves fast and accurate deduplication. Various experimental results highlight the practicality of FED for deduplication in large-scale datasets used for LLMs. We plan to make FED’s source code publicly available.
Limitations
As described in Section 3.1, unlike the baseline methods, FED performs comparisons for all pairs within a single bucket. While this strategy aims to reduce false negatives, it may result in excessive computations. Baseline methods achieve sufficient deduplication performance by selecting a representative document within each bucket and comparing it with the remaining ones. Thus, sensitivity experiments are required to determine how many comparisons are necessary for FED. If the same deduplication quality can be achieved with fewer comparisons, it would make FED faster. Detailed analysis and implementation are our future work.
Ethics Statement
We are committed to advancing NLP while adhering to ethical guidelines. Our study ensures compliance with data use policies, relying solely on publicly available datasets and open-source code, and strictly observing their licenses. Transparency is maintained through detailed descriptions of our methodologies, data sources, experimental processes, and the tools employed. Furthermore, we abstain from handling personal data or content that could perpetuate harm or discrimination against individuals or communities.
@@ Acknowledgement
Appendix
Figure 2: The process of MinHash generation.
Figure 3: Hashing by MinHash LSH.
Appendix A Jaccard Similarity
Jaccard similarity (Jaccard, 1912), $J(A,B)$, is a metric to quantify the similarity of two finite sets, $A$ and $B$:
$J(A,B)=\frac{|A\cap B|}{|A\cup B|}$ (6)
Appendix B MinHash Generation
Consider Figure 2 that illustrates the MinHash process. First, each document is broken down into smaller components, known as shingles or n-grams and represented by the set of its shingles. The shingles are typically sequences of contiguous words or characters. For example, the sentence "abcab" has a sequence of three shingles, $abc,bca,cab$. Figure 2 uses 3-grams for singling.
Then, we map the shingles to zero or positive integers using $H$ hash functions, $f_{1},f_{2},\cdots,f_{H}$. Each hash function assigns a unique number to a shingle, representing the document by a sequence of integers. As a result, we have $H$ sequences of integers for a document.
A MinHash signature for the document is a sequence of the minimum values obtained from each hash function, resulting in a sequence of length $H$, where $H$ is the total number of hash functions used. The Jaccard similarity between the two documents is then estimated by comparing their MinHash signatures. $H$ different hash functions result in a $H\times N$ signature matrix for $N$ documents, where each column represents the minimum hash values obtained from $H$ different hash functions.
Appendix C MinHash LSH
The critical difference between MinHash and MinHash LSH lies in the stage of generating duplicate pairs. In MinHash LSH, the signature column vector of each document is divided into $b$ bands, each of which has $r$ integers, as shown in Figure 3. LSH hashes the bands in each document to $k$ buckets. Let $S_{i}$ be the signature vector of the document $D_{i}$. For a given signature vector $S_{1}$ and $S_{2}$, when at least a pair of bands $B_{S_{1}}$ from $S_{1}$ and $B_{S_{2}}$ from $S_{2}$ hashed to the same bucket, we tag them as candidate pairs, potentially similar documents. For example, in Figure 3, $D_{2}$ and $D_{N}$ are potentially similar documents because they have bands hashed to the same bucket. For the candidate pairs, LSH performs pairwise comparisons on each document’s signature vectors of length $b\times r$. If the similarity between the two vectors exceeds a predefined threshold, the two documents are considered a duplicate pair. Once the union graph is constructed for each band, the union graphs for all bands are merged to obtain the final union graph.
Appendix D Suitability of FED’s Hash Functions
In general, a good hash function should have the following properties:
  • Determinacy: The same input produces the same output.
The proposed hash function satisfies these properties. Determinacy and uniformity are straightforward. Collision resistance is also met with a large value of $p$. When uniformity is satisfied, the probability of a collision approaches $\frac{1}{p}$, which is very small. In many contexts, hash functions are required to resist preimage attacks, making it computationally difficult to reverse-engineer the input from the hash value. However, it is not a concern because the hash function is not used for security purposes.
Appendix E Number of Buckets in LSH
For hashing to generate bucket IDs, FED sums the $r$ values in each band and uses the remainder of the division by the predefined constant number of buckets, $K$. This effectively maps each band, which is defined in the space $\mathbb{N}^{r}$, into the space $\mathbb{N}^{K}$. As $K$ increases, the process of scanning through files to group documents with the same key becomes more frequent. This process is proportional to $K$, resulting in $O(K)$. On the other hand, the number of documents with the same key is, on average, $O\left(\frac{N}{K}\right)$ when $N$ is the total number of documents. Since the comparison time is proportional to the square of the number of documents, the overall time complexity becomes $O\left(\left(\frac{N}{K}\right)^{2}K\right)=O\left(\frac{N^{2}}{K}\right)$. Thus, the processing time can be approximated as $T(K)=aNK+\frac{bN^{2}}{K}$, where $a$ and $b$ are some constants. To minimize $T(K)$, we differentiate it with respect to $K$ and set $\frac{dT(K)}{dK}=0$, yielding $K^{2}=\frac{bN^{2}}{aN}$. Solving for $K$, we find $K=\sqrt{\frac{b}{a}}N^{1/2}$. Thus, the optimal number of buckets can be expressed as $K=kN^{1/2}$, where $k=\sqrt{\frac{b}{a}}$ is a constant determined empirically (our experimental result shows that setting $k$ to a value between 1 and 2 is appropriate). Using this relationship, we identify the optimal value of $K$.

### v2

@@ 3.3 GPU-based Pairwise Comparison
Since typical GPU libraries do not support $sim_{sig}(S_{i},S_{j})$, we implement a custom GPU kernel. Replacing the equality operator (==absent$=$ =) with the multiplication operator ($\times$) makes the operation matrix multiplication. Thus, the GPU kernel performs $sim_{sig}(S_{i},S_{j})$ in a matrix multiplication-like manner and runs much faster than the ordinary kernel that implements straightforwardly $sim_{sig}(S_{i},S_{j})$. The kernel exploits tiling and shared memory Harris (2012) to optimize memory access patterns and enhance parallel processing performance. When computing the Jaccard similarity for all $c$ signature vectors in a pairwise manner, comparisons are performed only for pairs where $i<j$ to avoid redundant calculations. That is, instead of filling the entire $c$ x $c$ comparison result matrix, the computation is optimized to fill only the upper triangular matrix. Once the computation is complete, the index information of document pairs exceeding the predefined similarity threshold is sent to the CPU.
3.4 Parallel File I/O Operations
To perform pairwise comparisons of document signatures, we need to gather all documents that belong to the same bucket. That is, we must search all the stored files to locate all documents that belong to a given bucket. As Section 3.1 mentions, each process reads the entire set of hash result files and then extracts the document signatures corresponding to the $C$ bucket IDs, aggregating them into $C$ separate buffers. That is, each process handles $C$ bucket IDs concurrently, reducing the number of I/O operations. These buffers are then transferred to the GPU for pairwise comparisons. Each process iterates over buckets until all the buckets are handled. We choose $C$ such that (Total size of $\left(\frac{\text{Total size of files}}{\text{$K$}}\right)\times C\times N_{process}$ remains smaller than the available CPU memory, where $N_{process}$ is the number of processes at the CPU side in each node, and the total size of files is the sum of the sizes of all hash result files. $K$ is the number of buckets, ensuring that the aggregated data for the $C$ buckets fits within the system’s main memory. For example, when processing the RealNews dataset, we set $C=4096$. The number of buckets processed per file I/O operation $C$ is a tunable parameter. However, finding a suitable value that balances the CPU memory limits while sufficiently reducing the file I/O time is crucial.
Storing back hash results and bucket IDs
Before the pairwise comparison step of signatures, we transfer the hash results and bucket IDs to the CPU as shown in Figure 4(a). If the CPU memory is insufficient, we store them in secondary storage, such as SSDs and HDDs, to address the issues in the target environment with a limited main memory space.
3.5 Other Optimizations
Initial Processing by the CPU
The CPU processes handle different JSONL files in parallel. Each JSONL file is loaded into the CPU’s main memory using two buffers: one stores the text of all the documents in the file, and the other stores the document indices. After the buffers are filled with a batch of files, the buffer is transferred to the corresponding GPU for shingling and MinHash generation. The size of the buffers depends on the capacity of the GPU memory.
Shingling and MinHash generation on the GPU
The GPU performs shingling and generates MinHash values for the documents contained in the buffer. Each JSONL file consists of multiple documents, and each Streaming Multiprocessor (SM) Lindholm et al. (2008) within the GPU processes different documents independently. That is, a CUDA thread block processes a document, with each thread handling different MinHash functions.
Communication-computation overlapping
To maximize performance, we implement communication-computation overlapping techniques, such as double buffering Cheng et al. (2014). When the GPU performs computation, a buffer containing a batch of files is simultaneously transferred to the GPU. Transferring data generated by the GPU back to the CPU is also overlapped with GPU computation.
Hash parameters for GPU compatibility
The values for $p$ and $q$ described in Section 3.2 are carefully selected to address the limited support of the GPU’s 64-bit integer operations, ensuring that the result of the operation remains within the 32-bit integer range.
@@ 4 Experiment
Table 1: System configuration of the 4-node GPU cluster
GPU node
CPU 1 $\times$ AMD EPYC 7502 32-Core Processor
Memory 8 $\times$ 64GB DDR4 DIMM
GPU 4 $\times$ NVIDIA Tesla V100
OS Ubuntu 20.04.6 (kernel 5.4.0-100)
Compiler nvcc 12.4, GCC 9.4
GPU Driver 520.61.05
MPI Version 4.1
CUDA Version 12.4
Network Interface InfiniBand (200Gb/s)
Storage node
CPU
2 $\times$ AMD EPYC 7502 32-Core Processor
Memory
8 $\times$ 64GB DDR4 DIMM
RAID
Configuration
RAID5,
7 $\times$ 3.7TB Sabrent Rocket 4.0 Plus NVMe SSDs,
Max seq. read B/W per SSD: 7,000 MB/s,
Max seq. write B/W per SSD: 5,800 MB/s,
Max seq. read B/W of RAID5: 42, [^0] MB/s,
Max seq. write B/W of RAID5: 5,800 MB/s,
Total capacity: 22.2TB
File System
ext4
Network Interface
InfiniBand (200Gb/s)
4.1 Evaluation Environment
Target system configuration
We use a 4-node GPU cluster with a storage node for our experiments. Each node is equipped with four NVIDIA Tesla V100 GPUs, with each GPU having 32GB of memory. Table  1 provides the detailed target system configuration.
Table 2: The deduplication time in seconds on a single node is compared across the CPU baseline, the GPU baseline, and FED. The CPU baseline uses 64 logical CPU cores. Both the GPU baseline and FED are evaluated using four GPUs in a single node.
Method Dataset Generation Comparison Union Total
Speedup over
CPU baseline
Speedup over
GPU baseline
CPU baseline RealNews (100 GB) 18,361.1 2,983.7 17.6 21,361.4 1.0 -
(SlimPajama Shen et al. (2023)) Sampled C4 (100 GB) 22,748.4 3,878.8 33.1 26,660.3 1.0 -
GPU baseline RealNews (100 GB) 361.5 864.4 30.5 1,256.4 17.0 1.0
(NVIDIA NeMo Curator Jennings et al. (2024)) Sampled C4 (100 GB) 480.8 920.2 34.5 1,435.5 18.6 1.0
FED RealNews (100 GB) 70.6 126.4 2.4 199.5 107.2 6.3
Sampled C4 (100 GB) 91.7 243.9 2.6 338.2 78.8 4.2
Comparison baselines
As mentioned in Section 2.4, the CPU baseline is the CPU-based MinHash LSH implementation found in SlimPajama Shen et al. (2023). We identified and fixed some bugs in the implementation. The GPU baseline is the GPU-accelerated MinHash LSH implementation in NVIDIA NeMo Curator Jennings et al. (2024).
For MinHash LSH, we use the default settings of the CPU baseline, 128 hash functions with $b=16$ and $r=8$. Following the approach by Lee et al. Lee et al. (2022), we create shingles using five-grams and set the Jaccard similarity threshold to 0.8. The default settings of the CPU and GPU baselines are used for parameters other than the above-mentioned. As mentioned in Section 3.2, the number of buckets $K$ is obtained based on the number of documents $N$, such that $K=2N^{1/2}$.
Datasets used
The datasets we used in our experiments are the RealNews dataset Zellers et al. (2019) and C4 Raffel et al. (2020b). The RealNews dataset is a large English corpus of news articles from Common Crawl, and C4 is a filtered version of Common Crawl. For C4, we sampled 100GB out of the total 700GB. We choose these datasets for two reasons. They are large-scale datasets containing over 30 million documents. We select a sufficiently large dataset to understand how long the traditional MinHash LSH method takes to process it and to see FED achieves how much acceleration. It is known from a previous study Lee et al. (2022) that deduplication of the datasets we choose can improve the performance of language models.
Preprocessing the datasets
@@ 4.2 Processing Speed
We measure three different types of execution time for deduplication speed comparison: signature generation and bucketing, signature comparison, and union. The signature generation and bucketing time involves computing the MinHash signature matrices and assigning bucket IDs for each band. The signature comparison time involves finding documents with the same bucket ID and calculating the Jaccard similarity between them. Finally, the union time involves constructing a union graph structure for all documents and ultimately identifying the documents to be removed.
We used the complete 100GB RealNews dataset and the sampled C4 dataset, which consists of 100GB randomly sampled from the full 750GB C4 dataset, for this experiment. The RealNews and sampled C4 datasets contain 32M and 50M documents, respectively. These datasets are constructed to be the same size, allowing us to evaluate how FED performs on datasets of equal size but with different characteristics. The deduplication experiment results for the full C4 dataset can be found in subsection 4.4.
@@
When we measure the speed of the three deduplication methods, we flush page caches in the system to exclude their effect when performing file I/O. The CPU baseline processing time is measured using Python’s multiprocessing module across 64 logical CPU cores, while both the GPU baseline and FED are measured using four V100 GPUs in a single node. The results are summarized in Table 2. For deduplication on the RealNews dataset, FED is about 107.2 times faster than the CPU baseline and about 6.3 times faster than the GPU baseline. Notably, due to the introduction of reusable hash functions, the MinHash generation phase alone is roughly 260 times faster than the CPU baseline. Similarly, on the C4 dataset, FED achieves a speedup of 78.8 over the CPU baseline and about 4.2 over the GPU baseline.
@@
For the comparison stage, FED is about 23.6 times faster than the CPU baseline and 6.8 times faster than the GPU baseline on RealNews. Similarly, on the C4 dataset, FED achieves a speedup of 15.9 over the CPU baseline and 3.8 over the GPU baseline. Moreover, as described in Section 3.1, FED performs more signature comparisons than others. While it is possible to reduce the number of comparisons by following the method used in the CPU or GPU baseline, the FED implementation prioritizes minimizing false negatives by performing more comparisons. The result indicates that the GPU kernel optimization for pairwise signature comparison described in Section 3.5 is very effective, and in turn performing more comparisons is feasible because of the optimized GPU kernels.
@@
The processing speed of all methods is faster on the RealNews dataset than on the C4 dataset. It is attributed to the distinct characteristics of the two datasets. For both datasets, we use 100GB of data; however, RealNews consists of 32M documents with an average length of 3,335.17 characters (including spaces), whereas C4 consists of 50M documents with an average length of 2,191.15 characters (including spaces). Consequently, the higher number of documents in C4 increases the time taken for the generation and comparison phases.
Table 3: Comparing the deduplication accuracy of the MinHash LSH-based approaches and the standard MinHash algorithm.
0.1M documents 1M documents
Ratio Jaccard Ratio Jaccard
Standard MinHash 7,317 / 0.1M - 289,229 / 1M -
CPU baseline 7,284 / 0.1M 0.995 288,683 / 1M 0.998
GPU baseline 7,144 / 0.1M 0.953 285,134 / 1M 0.976
FED 7,175 / 0.1M 0.956 283,248 / 1M 0.966
@@ 4.3 Deduplication Accuracy
We use a 0.1M and 1M dataset sampled from RealNews for this experiment because the original MinHash algorithm is time-consuming on large datasets. Our sampling is not entirely random to ensure many duplicated pairs even within the two smaller datasets. We first use MinHash LSH to identify similar groups, then select a proportion of each group to construct the subset.
@@
We evaluate the accuracy of the deduplication methods by measuring the Jaccard similarity between the set of all near duplicates obtained using the original and standard MinHash and that obtained using each method. A higher similarity indicates that the method identifies duplicate pairs more similarly to the standard MinHash.
Table 3 compares the deduplication accuracy of each method. The Ratio column represents the total number of documents that have near-duplicate pairs divided by the total number of documents. Jaccard column shows the Jaccard similarity between the sets of all near duplicates by the standard MinHash and each method. FED achieves a Jaccard similarity of 0.95 or higher to the standard MinHash result. A similarity score of 0.95 or higher means that the method has identified near-duplicate documents almost identically to the standard MinHash. In fact, on the 1M dataset, out of the 283,248 documents identified as having duplicate pairs by FED, 281,355 overlapped with those identified by MinHash. This indicates that FED not only achieves faster deduplication than other methods but also identifies near-duplicate documents similar to the standard MinHash.
@@
Generation 92.9 81.3 70.6 53.6 46.9
Comparison 377.1 216.9 126.4 92.0 62.2
Union 1.5 2.0 2.4 2.4 2.6
Total Time 471.5 300.2 199.4 148.0 111.7
@@ 4.4 Scalability
Table 6 shows the results of measuring execution time when the number of GPUs is varying. Note that eight or more GPUs indicate a multi-node system configuration. We use the RealNews dataset, which has 32 million documents. Even with just a single GPU, the method is 45 times faster than the CPU baseline, and with 16 GPUs, the deduplication process completes in less than 2 minutes. This experiment demonstrates that our method works effectively in multi-node environments, and with more available node resources, even faster processing is possible.
@@
We increase the dataset size to verify whether FED can process trillions of tokens at a practical processing speed. The results of deduplication using 16 GPUs on the RealNews and RedPajama-1T Computer (2023) datasets are shown in Table 6. It shows that FED can process trillions of tokens in just 6 hours. Since recent LLMs are typically trained with more than trillions of tokens, the results indicate that FED can be practically used to preprocess the training data for LLMs.
@@
The number of hash functions used in the MinHash generation step is a key parameter that could impact the deduplication performance. For example, Rae et al. (2021b) use 450 hash functions. We vary the number of hash functions from 128 to 256 and 512, and Table 6 shows the results. Despite changing the number of hash functions, FED maintained a Jaccard similarity of 0.92 or higher to the standard MinHash results, demonstrating that FED remains robust even with variations in the number of hash functions.
@@ 5 Conclusion
As the amount of data used for training LLMs continues to grow, the time required for data preprocessing has become increasingly significant. We present a framework called FED, which performs MinHash LSH-based deduplication efficiently on GPUs and uses novel hash functions. FED significantly outperforms the CPU-based deduplication tool (using 64 logical CPU cores) included in SlimPajama by up to 107.2 times and the GPU-based deduplication tool included in NVIDIA NeMo Curator by up to 6.3 times with a node of four GPUs when processing a dataset with 30 billion tokens. In a multi-node and multi-GPU environment, deduplication of 30 billion tokens by FED takes about 111.67 seconds and 1.2 trillion tokens about six hours. Similar documents identified by FED show a Jaccard similarity of 0.95 or higher with those found using the standard MinHash algorithm, indicating that FED achieves fast and accurate deduplication. Various experimental results highlight the practicality of FED for deduplication in large-scale datasets used for LLMs. We also make FED’s source code publicly available.